# Transmission Conditions for the Non-Overlapping Schwarz Coupling of Full Order and Operator Inference Models

### v1

@@ 3.2 Numerical results
Our first set of numerical results examines the influence of the Robin boundary condition parameters, $\alpha_{ij}$ and $\beta_{ij}$, as defined in Section 3.1, on: (i) the error between the converged Schwarz solution and the corresponding monolithic solution, and (ii) the number of Schwarz iterations required to satisfy the prescribed convergence tolerance. This is achieved by performing a parameter sweep for all four $\alpha_{ij}$ and $\beta_{ij}$ parameters within the 1D wave propagation problem setting described in Section 2.1. The parameter spaces used in the sweep are:
@@
where $\sigma_{\max}=\max_{t,\,x}\sigma(x,t)$ denotes the maximum stress within the bar for the corresponding monolithic solution. Normalizing the $\alpha_{ij}$ parameter by a reference stress value compensates for the large difference in orders of magnitude between the displacements and tractions used in the Robin transmission conditions and improves the interpretability of the results. The parameter spaces defined in (12) were selected to capture the full range of variation in both the error and number of iterations required for convergence.
The error between the converged Schwarz solution and the corresponding monolithic solution is calculated at each time step. For subdomain $k$ at time $t_{n}$ we define:
$\varepsilon_{k}(t_{n})=\big{\lVert}\Delta~u_{k}(t_{n})\big{\rVert}_{2}+\Delta t\,\big{\lVert}\Delta~v_{k}(t_{n})\big{\rVert}_{2}+\tfrac{1}{2}\,\Delta t^{2}\,\big{\lVert}\Delta~a_{k}(t_{n})\big{\rVert}_{2}$ (13)
@@
$\Delta~u_{k}(t_{n})$ $=~u_{k}^{\mathrm{SCH}}(t_{n})-~u_{k}^{\mathrm{MON}}(t_{n})$ (14)
$\Delta~v_{k}(t_{n})$ $=~v_{k}^{\mathrm{SCH}}(t_{n})-~v_{k}^{\mathrm{MON}}(t_{n})$
$\Delta~a_{k}(t_{n})$ $=~a_{k}^{\mathrm{SCH}}(t_{n})-~a_{k}^{\mathrm{MON}}(t_{n})$
and $\lVert\cdot\rVert_{2}$ denotes the Euclidean norm. Here $u^{\mathrm{SCH}}$, $v^{\mathrm{SCH}}$, $a^{\mathrm{SCH}}$ are the converged Schwarz displacement, velocity and acceleration, and $u^{\mathrm{MON}}$, $v^{\mathrm{MON}}$, $a^{\mathrm{MON}}$ are the corresponding monolithic quantities. Lastly, the time-averaged error is calculated as:
$\varepsilon_{avg}=\frac{1}{N_{t}-1}\sum_{i=1}^{N_{k}}\sum_{n=2}^{N_{t}}\varepsilon_{k}(t_{n})$ (15)
@@
A Pareto plot of the parameter sweep results is presented in Figure 2. The error, as defined in (15), is plotted against the mean number of Schwarz iterations required for convergence for varying $\alpha_{ij}$ and $\beta_{ij}$ parameters. The color mapping represents the ratio $\beta_{12}/\overline{\alpha}_{12}$, serving as an indicator of the relative magnitude of Dirichlet to Neumann data used in the Robin transmission condition. Here, $\overline{\alpha}_{ij}=\sigma_{max}\,\cdot\alpha_{ij}$. A high ratio (red) approaches a pure Dirichlet condition, whereas a low ratio (blue) approaches a pure Neumann condition. For brevity, the plot showing results for $\beta_{21}/\overline{\alpha}_{21}$ is omitted, as similar trends are observed. The result obtained with an alternating Dirichlet–Neumann transmission condition is also included in gray as a reference. No relaxation was employed in generating our alternating Dirichlet-Neumann results, as numerical experiments revealed that the inclusion of relaxation worsened both accuracy and efficiency of the NO-SAM algorithm.
Fig. 2: FOM-FOM coupling: Pareto plot illustrating the relationship between the error in the converged Schwarz solution relative to the corresponding monolithic solution and the mean number of Schwarz iterations required for convergence, for varying $\alpha_{ij}$ and $\beta_{ij}$ parameters. The unrelaxed alternating Dirichlet-Neumann result is also plotted as a reference in gray.
@@
Fig. 3: FOM-FOM coupling: plots of the NO-SAM displacement, velocity, and acceleration solutions obtained with alternating Dirichlet–Neumann transmission conditions, compared to the corresponding monolithic solution (dashed line). The blue subdomain employs a Dirichlet transmission condition, and the red subdomain employs a Neumann transmission condition.
@@
Fig. 4: FOM-FOM coupling: plots of the NO-SAM displacement, velocity, and acceleration solutions obtained with Robin–Robin transmission conditions (lowest error parameter set, $\left[\overline{\alpha}_{12},\overline{\alpha}_{21},\beta_{12},\beta_{21}\right]=\left[10^{-3},10^{-3},1,1\right]$) compared to the corresponding monolithic solution (dashed line). Solutions computed using NO-SAM in $\Omega_{1}$ and $\Omega_{2}$ are shown in blue and red, respectively.
@@
Fig. 5: FOM-FOM coupling: plots of the NO-SAM displacement, velocity, and acceleration solutions obtained with Robin–Robin transmission conditions (highest error parameter set, $\left[\overline{\alpha}_{12},\overline{\alpha}_{21},\beta_{12},\beta_{21}\right]=\left[10^{-1},1,1,3\right]$) compared to the corresponding monolithic solution (dashed line). Solutions computed using NO-SAM in $\Omega_{1}$ and $\Omega_{2}$ are shown in blue and red, respectively.
@@ 4.1 OpInf ROM overview
$U=\left[~u(t_{1}),\,~u(t_{2}),\,\dots\,,~u(t_{\tau})\right]\in\mathbb{R}^{N\times\tau}$ (16)
It is important to note that snapshot velocities $\dot{~u}(t_{n})$ and accelerations $\ddot{~u}(t_{n})$ can also be included in the snapshot matrix U~U along with the displacements, though this is not considered herein. The reduced basis is obtained by performing the singular value decomposition of $U=~\Phi~\Sigma~V^{T}$, where $\Phi\in\mathbb{R}^{N\times N}$ is an orthogonal matrix whose first $R=\operatorname{rank}(~U)>0$ columns span the column space of U~U. The first $r\leq R$ columns of $\Phi$ define an optimal $r$-dimensional basis for the columns of U~U in the $l^{2}$ sense. We denote this truncated basis by $\Phi_{r}\in\mathbb{R}^{N\times r}$. The fraction of statistical energy captured by a basis of size $r$ is defined as
$E(r)\coloneq\frac{\sum_{i=1}^{r}\sigma_{i}^{2}}{\sum_{i=1}^{R}\sigma_{i}^{2}},$ (17)
where $\sigma_{i}$ denotes the $i^{th}$ singular value of $\Sigma$ (not to be confused with the stress notation used previously).
@@
For standard intrusive and projection-based ROMs, reduced operators are obtained by projecting the full order operators discretizing the governing PDEs onto the span of the reduced basis using $\Phi$. The main disadvantage of this procedure is that it requires explicit access to the full order operators defining the governing problem. One selling point of OpInf ROMs is that the reduced operators are non-intrusively inferred directly from the snapshot data used to build $\Phi$, eliminating the need to access the full order system operators.
@@
The first step in developing an appropriate OpInf ROM for our model problem (Section 2) involves writing the functional form corresponding to the FOM semi-discretization. It is straight-forward to see that, for a linear elastic material and no body force, discretizing (3) in space using the finite element method results in the following functional form:
$M\ddot{~u}+~K~u=~F,$ (18)
where
$F_{i}:=\int_{\partial_{T}\Omega}~T\cdot\boldsymbol{\xi}_{i}\,dS$ (19)
@@
for the $i^{\mathrm{th}}$ test function $\boldsymbol{\xi}_{i}$. Our goal is to derive the functional form of a projection-based ROM corresponding to (18). Suppose the prescribed traction T:=t~T:=~t is piecewise constant on the element faces along the subdomain interface $\Gamma$. In this case, (19) can be approximated as $F_{i}=~t\cdot\int_{\Gamma}\boldsymbol{\xi}_{i}dS$, so that F~F takes the form: F=Ht~F=~H~t for a given prescribed traction vector, t~t. The resulting discrete system $M\ddot{~u}+~K~u=~H~t$ incorporates the traction, but not displacement boundary conditions. Let $i_{d}$ denote the solution indices on which the displacement boundary conditions are imposed, and let $i_{u}$ denote the remaining unconstrained indices. Suppose g~g denotes the vector of displacement values imposed at $i_{d}$. Then, (18) can be rewritten in the form
$\bar{~M}\,\ddot{\bar{~u}}+\bar{~K}\,\bar{~u}=\bar{~H}\,~t+\bar{~B}\,~g,$ (20)
where $\bar{~M}:=~M(~i_{u},~i_{u})$, $\bar{~K}:=~K(~i_{u},~i_{u})$, $\bar{~u}:=~u(~i_{u})$, $\bar{~H}:=~H(~i_{u},:)$ and $\bar{~B}:=~K(~i_{u},~i_{d})$.
@@
$\hat{~M}\,\ddot{\hat{~u}}+\hat{~K}\,\hat{~u}=\hat{~H}\,~t+\hat{~B}\,~g,$ (21)
@@
where the “hat” notation denotes reduced order operators. To learn the reduced order operators $\hat{~M},\hat{~K}\in\mathbb{R}^{r\times r}$, $\hat{~H}\in\mathbb{R}^{r\times s}$, and $\hat{~B}\in\mathbb{R}^{r\times m}$ in the monolithic OpInf setting, we solve the following linear regression problem
$\min_{\hat{~M},\hat{~K},\hat{~H},\hat{~B}}\sum_{p=1}^{j}\lVert\hat{~M}\,\ddot{\hat{~u}}_{p}+\hat{~K}\,\hat{~u}_{p}-\hat{~H}\,~t-\hat{~B}\,~g,\rVert^{2}_{2}$ (22)
@@
given boundary conditions $t\in\mathbb{R}^{s}$ and $g\in\mathbb{R}^{m}$. Here, $\hat{~u}_{p}=~\Phi_{r}^{T}~u(t_{p})$ and $\ddot{\hat{~u}}_{p}=~\Phi_{r}^{T}\ddot{~u}(t_{p})$ are obtained from the displacement and acceleration snapshot matrices, respectively, and $j\leq\tau$ denotes the number of snapshots used from the FOM dataset. If the acceleration snapshots are not readily available, they can be approximated using a finite difference scheme. To avoid the trivial solution while solving (22) in a fully unconstrained fashion—where each reduced operator identically equal to zero constitutes a valid minimum—we instead solve a modified linear regression problem obtained by applying the inverse of the reduced mass matrix to each term in (21)
$\min_{\tilde{~K},\tilde{~H},\tilde{~B}}\sum_{p=1}^{j}\lVert\ddot{\hat{~u}}_{p}+\tilde{~K}\,\hat{~u}_{p}-\tilde{~H}\,~t-\tilde{~B}\,~g,\rVert^{2}_{2}$ (23)
@@
where the tilde notation denotes operators that have been premultiplied by the inverse of the reduced mass matrix, i.e. $\tilde{~K}=\hat{~M}^{-1}\hat{~K}$, $\tilde{~H}=\hat{~M}^{-1}\hat{~H}$, and $\tilde{~B}=\hat{~M}^{-1}\hat{~B}$. This formulation is employed in the numerical examples presented below. Once the reduced operators are determined, boundary conditions in the OpInf context are imposed directly by specifying g~g and t~t in (21).
@@
Remark 4.2. While the assumption of constant traction at the interface is trivial in the 1D setting considered here, extending this formulation to higher dimensions may introduce additional challenges, and require alternate formulations/implementations. One possible alternate formulation assumes that the quadrature leading to F~F in (19) can be written as $F=~C~t^{\prime}$ for some constant matrix C~C and a known (possibly nonlinear) function $t^{\prime}=~F(~t)$. In this case, it is possible to learn reduced versions of C~C to construct the ROM. In particular, using $\tilde{~u}=\boldsymbol{\Phi}_{r}\hat{~u}$ leads to
$\hat{~M}\,\ddot{\hat{~u}}+\hat{~K}\,\hat{~u}=\hat{~C}\,~t^{\prime}+\hat{~B}\,~g,$ (24)
@@
where $\hat{~C}=\boldsymbol{\Phi}_{r}^{T}~C$. Learning the operators in (24) is expected to lead to a viable ROM for each t~t. We plan to explore this research direction in future work.
@@
$\min_{\tilde{~K},\tilde{~H},\tilde{~B}}\sum_{p=1}^{j}\Big{\lVert}\ddot{\hat{~u}}_{p}+\tilde{~K}\,\hat{~u}_{p}-\tilde{~H}\,~t-\tilde{~B}\,~g\Big{\rVert}_{2}^{2}+\lambda^{2}\Big{(}\lVert\tilde{~K}\rVert_{F}^{2}+\lVert\tilde{~H}\rVert_{F}^{2}+\lVert\tilde{~B}\rVert_{F}^{2}\Big{)}$ (25)

### v2

@@ 3.2 Numerical results
Our first set of numerical results examines the influence of the Robin boundary condition parameters, $\alpha_{ij}$ and $\beta_{ij}$, as defined in Section 3.1, on: (i) the error between the converged Schwarz solution and the corresponding monolithic solution, and (ii) the number of Schwarz iterations per time step required to satisfy the prescribed convergence tolerance. This is achieved by performing a parameter sweep for all four $\alpha_{ij}$ and $\beta_{ij}$ parameters within the 1D wave propagation problem setting described in Section 2.1. The parameter spaces used in the sweep are:
@@
where $\sigma_{\max}:=\max_{t,\,x}\sigma(x,t)$ denotes the maximum stress within the bar for the corresponding monolithic solution. Normalizing the $\alpha_{ij}$ parameter by a reference stress value compensates for the large difference in orders of magnitude between the displacements and tractions used in the Robin transmission conditions and improves the interpretability of the results. The parameter spaces defined in (12) were selected to capture the full range of variation in both the error and number of iterations required for convergence.
The error between the converged Schwarz solution and the corresponding monolithic solution is calculated at each time step. For subdomain $k$ at time $t_{n}$ we define
$\varepsilon_{k}(t_{n}):=\big\lVert\Delta\mathord{\btensor u}_{k}(t_{n})\big\rVert_{2}+\Delta t\,\big\lVert\Delta\mathord{\btensor v}_{k}(t_{n})\big\rVert_{2}+\tfrac{1}{2}\,\Delta t^{2}\,\big\lVert\Delta\mathord{\btensor a}_{k}(t_{n})\big\rVert_{2}$ (13)
@@
$\Delta\mathord{\btensor u}_{k}(t_{n})$ $=\mathord{\btensor u}_{k}^{\mathrm{SCH}}(t_{n})-\mathord{\btensor u}_{k}^{\mathrm{MON}}(t_{n})$ (14)
$\Delta\mathord{\btensor v}_{k}(t_{n})$ $=\mathord{\btensor v}_{k}^{\mathrm{SCH}}(t_{n})-\mathord{\btensor v}_{k}^{\mathrm{MON}}(t_{n})$
$\Delta\mathord{\btensor a}_{k}(t_{n})$ $=\mathord{\btensor a}_{k}^{\mathrm{SCH}}(t_{n})-\mathord{\btensor a}_{k}^{\mathrm{MON}}(t_{n})$
and $\lVert\cdot\rVert_{2}$ denotes the Euclidean norm. Here $\mathord{\btensor u}^{\mathrm{SCH}}$, $\mathord{\btensor v}^{\mathrm{SCH}}$, $\mathord{\btensor a}^{\mathrm{SCH}}$ are the converged Schwarz displacement, velocity and acceleration, and $\mathord{\btensor u}^{\mathrm{MON}}$, $\mathord{\btensor v}^{\mathrm{MON}}$, $\mathord{\btensor a}^{\mathrm{MON}}$ are the corresponding monolithic quantities. Lastly, the time-averaged error is defined as
$\varepsilon_{avg}:=\frac{1}{N_{t}-1}\sum_{k=1}^{N_{k}}\sum_{n=2}^{N_{t}}\varepsilon_{k}(t_{n})$ (15)
@@
A Pareto plot of the parameter sweep results is presented in Figure 2. The error, as defined in (15), is plotted against the mean number of Schwarz iterations per time step required for convergence for varying $\alpha_{ij}$ and $\beta_{ij}$ parameters. The color mapping represents the ratio $\beta_{12}/\overline{\alpha}_{12}$, serving as an indicator of the relative magnitude of Dirichlet to Neumann data used in the Robin transmission condition. Here, $\overline{\alpha}_{ij}:=\sigma_{max}\,\cdot\alpha_{ij}$. A high ratio (red) approaches a pure Dirichlet condition, whereas a low ratio (blue) approaches a pure Neumann condition. It is curious that the near Dirichlet-Dirichlet case gives the most accurate and efficient results, given that Dirichlet-Dirichlet NO-SAM is known not to converge [28]. This behavior was also observed in our implementation. For brevity, the plot showing results for $\beta_{21}/\overline{\alpha}_{21}$ is omitted, as similar trends are observed. The result obtained with an alternating Dirichlet–Neumann transmission condition is also included in gray as a reference. No relaxation was employed in generating our alternating Dirichlet-Neumann results, as numerical experiments revealed that the inclusion of relaxation worsened both accuracy and efficiency of the NO-SAM algorithm.
Fig. 2: FOM-FOM coupling: Pareto plot illustrating the relationship between the error in the converged Schwarz solution relative to the corresponding monolithic solution and the mean number of Schwarz iterations per time step required for convergence, for varying $\alpha_{ij}$ and $\beta_{ij}$ parameters. The unrelaxed alternating Dirichlet-Neumann result is also plotted as a reference in gray.
@@
Fig. 3: FOM-FOM coupling: plots of the NO-SAM displacement, velocity, and acceleration solutions obtained with alternating Dirichlet–Neumann transmission conditions, compared to the corresponding monolithic solution (dashed line). The blue subdomain $\Omega_{1}$ employs a Dirichlet transmission condition, and the red subdomain $\Omega_{2}$ employs a Neumann transmission condition.
@@
Fig. 4: FOM-FOM coupling: plots of the NO-SAM displacement, velocity, and acceleration solutions obtained with Robin–Robin transmission conditions (lowest error parameter set, $\left[\overline{\alpha}_{12},\overline{\alpha}_{21},\beta_{12},\beta_{21}\right]=\left[10^{-3},10^{-3},1,1\right]$) compared to the corresponding monolithic solution (dashed line). The solutions computed for subdomains $\Omega_{1}$ and $\Omega_{2}$ are shown in blue and red, respectively.
@@
Fig. 5: FOM-FOM coupling: plots of the NO-SAM displacement, velocity, and acceleration solutions obtained with Robin–Robin transmission conditions (highest error parameter set, $\left[\overline{\alpha}_{12},\overline{\alpha}_{21},\beta_{12},\beta_{21}\right]=\left[10^{-1},1,1,3\right]$) compared to the corresponding monolithic solution (dashed line). The solutions computed for subdomains $\Omega_{1}$ and $\Omega_{2}$ are shown in blue and red, respectively.
@@ 4.1 Monolithic OpInf ROM overview
$\mathord{\btensor U}:=\left[\mathord{\btensor u}(t_{1}),\,\mathord{\btensor u}(t_{2}),\,\dots\,,\mathord{\btensor u}(t_{\tau})\right]\in\mathbb{R}^{N\times\tau}$ (16)
[…]
$E(r)\coloneq\frac{\sum_{i=1}^{r}\mu_{i}^{2}}{\sum_{i=1}^{R}\mu_{i}^{2}}$ (17)
where $\mu_{i}$ denotes the $i^{th}$ singular value of .
@@
For standard intrusive and projection-based ROMs, reduced operators are obtained by projecting the full order operators discretizing the governing PDEs onto the span of the reduced basis using . The main disadvantage of this procedure is that it requires explicit access to the full order operators defining the governing problem. One selling point of OpInf ROMs is that the reduced operators are non-intrusively inferred directly from the snapshot data used to build , eliminating the need to access the full order system operators.
@@
The first step in developing an appropriate OpInf ROM for our model problem (Section 2.1) is to establish the functional form of the FOM semi-discretization. As outlined in Section 2, applying the finite element method to discretize the weak form (3) in space yields the functional representation given in (4). In the absence of body forces, the force vector reduces to
$F_{i}=\int_{\partial_{T}\Omega}\mathord{\btensor T}\cdot\boldsymbol{\xi}_{i}\ dS$ (18)
@@
for the $i^{\mathrm{th}}$ test function $\boldsymbol{\xi}_{i}$. Our goal is to derive the functional form of a projection-based ROM corresponding to (4). Suppose the prescribed traction $\mathord{\btensor T}:=\mathord{\btensor t}$ is piecewise constant on the element faces along the subdomain interface $\Gamma$. In this case, (18) can be approximated as $F_{i}=\mathord{\btensor t}\cdot\int_{\Gamma}\boldsymbol{\xi}_{i}dS$, so that $\mathord{\btensor F}$ takes the form: $\mathord{\btensor F}=\mathord{\btensor H}\mathord{\btensor t}$ for a given prescribed traction vector, $\mathord{\btensor t}$. The resulting discrete system $\mathord{\btensor M}\ddot{\mathord{\btensor u}}+\mathord{\btensor K}\mathord{\btensor u}=\mathord{\btensor H}\mathord{\btensor t}$ incorporates the traction, but not displacement boundary conditions. Let $\mathord{\btensor i}_{d}$ denote the solution indices on which the displacement boundary conditions are imposed, and let $\mathord{\btensor i}_{u}$ denote the remaining unconstrained indices. Suppose $\mathord{\btensor g}$ denotes the vector of displacement values imposed at $\mathord{\btensor i}_{d}$. Then, (4) can be rewritten in the form
$\bar{\mathord{\btensor M}}\,\ddot{\bar{\mathord{\btensor u}}}+\bar{\mathord{\btensor K}}\,\bar{\mathord{\btensor u}}=\bar{\mathord{\btensor H}}\,\mathord{\btensor t}+\bar{\mathord{\btensor B}}\,\mathord{\btensor g}$ (19)
where $\bar{\mathord{\btensor M}}:=\mathord{\btensor M}\left[\mathord{\btensor i}_{u},\mathord{\btensor i}_{u}\right]$, $\bar{\mathord{\btensor K}}:=\mathord{\btensor K}\left[\mathord{\btensor i}_{u},\mathord{\btensor i}_{u}\right]$, $\bar{\mathord{\btensor u}}:=\mathord{\btensor u}\left[\mathord{\btensor i}_{u}\right]$, $\bar{\mathord{\btensor H}}:=\mathord{\btensor H}\left[\mathord{\btensor i}_{u},:\right]$ and $\bar{\mathord{\btensor B}}:=-\mathord{\btensor K}\left[\mathord{\btensor i}_{u},\mathord{\btensor i}_{d}\right]$.
@@
$\hat{\mathord{\btensor M}}\,\ddot{\hat{\mathord{\btensor u}}}+\hat{\mathord{\btensor K}}\,\hat{\mathord{\btensor u}}=\hat{\mathord{\btensor H}}\,\mathord{\btensor t}+\hat{\mathord{\btensor B}}\,\mathord{\btensor g}$ (20)
@@
where the “hat” notation denotes reduced order operators. To learn the reduced order operators $\hat{\mathord{\btensor M}},\hat{\mathord{\btensor K}}\in\mathbb{R}^{r\times r}$, $\hat{\mathord{\btensor H}}\in\mathbb{R}^{r\times l}$, and $\hat{\mathord{\btensor B}}\in\mathbb{R}^{r\times m}$ in the monolithic OpInf setting, we solve the following linear regression problem
$\min_{\hat{\mathord{\btensor M}},\hat{\mathord{\btensor K}},\hat{\mathord{\btensor H}},\hat{\mathord{\btensor B}}}\sum_{p=1}^{P}\lVert\hat{\mathord{\btensor M}}\,\ddot{\hat{\mathord{\btensor u}}}(t_{p})+\hat{\mathord{\btensor K}}\,\hat{\mathord{\btensor u}}(t_{p})-\hat{\mathord{\btensor H}}\,\mathord{\btensor t}(t_{p})-\hat{\mathord{\btensor B}}\,\mathord{\btensor g}(t_{p})\rVert^{2}_{2}$ (21)
@@
given boundary condition data $\mathord{\btensor t}(t_{p})\in\mathbb{R}^{l}$ and $\mathord{\btensor g}(t_{p})\in\mathbb{R}^{m}$. These terms are expressed as functions of time to account for cases where they are not constant. In such instances, snapshot data of the boundary conditions on the domain boundary $\partial\Omega$ must be collected at each time step. The reduced-order quantities $\hat{\mathord{\btensor u}}(t_{p}):={}_{r}^{T}\mathord{\btensor u}(t_{p})$ and $\ddot{\hat{\mathord{\btensor u}}}(t_{p}):={}_{r}^{T}\ddot{\mathord{\btensor u}}(t_{p})$ are obtained from the displacement and acceleration snapshot matrices, respectively, where $P\leq\tau$ denotes the total number of snapshots taken from the FOM dataset. If acceleration snapshots are not available, they may be approximated using a finite-difference scheme. To avoid the trivial solution while solving (21) in a fully unconstrained fashion—where each reduced operator identically equal to zero constitutes a valid minimum—we instead solve a modified linear regression problem obtained by applying the inverse of the reduced mass matrix to each term in (20)
$\min_{\tilde{\mathord{\btensor K}},\tilde{\mathord{\btensor H}},\tilde{\mathord{\btensor B}}}\sum_{p=1}^{P}\lVert\ddot{\hat{\mathord{\btensor u}}}(t_{p})+\tilde{\mathord{\btensor K}}\,\hat{\mathord{\btensor u}}(t_{p})-\tilde{\mathord{\btensor H}}\,\mathord{\btensor t}(t_{p})-\tilde{\mathord{\btensor B}}\,\mathord{\btensor g}(t_{p})\rVert^{2}_{2}$ (22)
@@
where the tilde notation denotes operators that have been premultiplied by the inverse of the reduced mass matrix, i.e. $\tilde{\mathord{\btensor K}}:=\hat{\mathord{\btensor M}}^{-1}\hat{\mathord{\btensor K}}$, $\tilde{\mathord{\btensor H}}:=\hat{\mathord{\btensor M}}^{-1}\hat{\mathord{\btensor H}}$, and $\tilde{\mathord{\btensor B}}:=\hat{\mathord{\btensor M}}^{-1}\hat{\mathord{\btensor B}}$. This formulation is employed in the numerical examples presented below. Once the reduced operators are determined, boundary conditions in the OpInf context are imposed directly by specifying $\mathord{\btensor g}$ and $\mathord{\btensor t}$ in (20).
@@
Remark 4.2. While the assumption of constant traction at the interface is trivial in the 1D setting considered here, extending this formulation to higher dimensions may introduce additional challenges, and require alternate formulations/implementations. One possible alternate formulation assumes that the quadrature leading to $\mathord{\btensor F}$ in (18) can be written as $\mathord{\btensor F}=\mathord{\btensor C}\mathord{\btensor t}^{\prime}$ for some constant matrix $\mathord{\btensor C}$ and a known (possibly nonlinear) function $\mathord{\btensor t}^{\prime}=\mathord{\btensor F}(\mathord{\btensor t})$. In this case, it is possible to learn reduced versions of $\mathord{\btensor C}$ to construct the ROM. In particular, we have
$\hat{\mathord{\btensor M}}\,\ddot{\hat{\mathord{\btensor u}}}+\hat{\mathord{\btensor K}}\,\hat{\mathord{\btensor u}}=\hat{\mathord{\btensor C}}\,\mathord{\btensor t}^{\prime}+\hat{\mathord{\btensor B}}\,\mathord{\btensor g}$ (23)
@@
where $\hat{\mathord{\btensor C}}:=\boldsymbol{\Phi}_{r}^{T}\mathord{\btensor C}$. Learning the operators in (23) is expected to lead to a viable ROM for each $\mathord{\btensor t}$. We plan to explore this research direction in future work.
@@
$\min_{\tilde{\mathord{\btensor K}},\tilde{\mathord{\btensor H}},\tilde{\mathord{\btensor B}}}\sum_{p=1}^{P}\Big\lVert\ddot{\hat{\mathord{\btensor u}}}(t_{p})+\tilde{\mathord{\btensor K}}\,\hat{\mathord{\btensor u}}(t_{p})-\tilde{\mathord{\btensor H}}\,\mathord{\btensor t}(t_{p})-\tilde{\mathord{\btensor B}}\,\mathord{\btensor g}(t_{p})\Big\rVert_{2}^{2}+\lambda^{2}\Big(\lVert\tilde{\mathord{\btensor K}}\rVert_{F}^{2}+\lVert\tilde{\mathord{\btensor H}}\rVert_{F}^{2}+\lVert\tilde{\mathord{\btensor B}}\rVert_{F}^{2}\Big)$ (24)